Show that the plane 3x+12y=62-17 = 0, touches the conicoid 3x
1
Expert's answer
2021-09-21T02:16:05-0400

Show that the plane "3x+12y-6z-17 = 0," touches the conicoid "3x^2-6y^2+9z^2+17=0" and find the point of contact.



"F(x, y, z)=3x^2-6y^2+9z^2+17=0"

"F_x=6x, F_y=-12y, F_z=18z"

The equation of the tangent plane is


"F_x(x-x_0)+F_y(y-y_0)+F_z(z-z_0)=0"

Substitute


"6x_0(x-x_0)-12y_0(y-y_0)+18z_0(z-z_0)=0"

"3x_0(x-x_0)-6y_0(y-y_0)+9z_0(z-z_0)=0"

"3x_0x-6y_0y+9z_0z-(3x_0^2-6y_0^2+9z_0^2)=0"

"3x_0x-6y_0y+9z_0z+17=0"

"-3x_0x+6y_0y-9z_0z-17=0"

Given the plane "3x+12y-6z-17 = 0." Then


"x_0=-1, y_0=2,z_0=2\/3"

The plane "3x+12y-6z-17 = 0," touches the conicoid "3x^2-6y^2+9z^2+17=0" at the point "(-1,2,2\/3)."